The height of a triangle is 5 more than the base. If the area of the triangle is 18 square units then how long is the base of the triangle?
Answer by checkley79(3341)

area=bh/2
h=b+5
18=b(b+5)/2
18=(b^2+5b)/2 cross multiply.
b^2+5b=18*2
b^2+5b=36
b^2+5b-36=0
(b+9)(b-4)=0
b-4=0
b=4 ans. for the base.
h=4+5=9 ans. for the height.
Proof:
18=4*9/2
18=36/2
18=18